OpenStudy (anonymous):

The legs of a right triangle are 4 units and 7 units. What is the length of the hypotenuse?

OpenStudy (alexandervonhumboldt2):

4^2+7^2=c^2

OpenStudy (alexandervonhumboldt2):

16+49=c^2

OpenStudy (alexandervonhumboldt2):

c=sqrt65

OpenStudy (anonymous):

To find the answer you do 4^2*7^2= you answer squared

OpenStudy (alexandervonhumboldt2):

cannot be simplified i think

OpenStudy (anonymous):

@AlexandervonHumboldt2 thanks!

OpenStudy (alexandervonhumboldt2):

np